OpenStudy (anonymous):

lien earned $1,200 more than her husband. together they earned $42,900. how much did each of them earn?

OpenStudy (mertsj):

x = amount her husband earned x+1200= amount lien earned x+(x+1200)=42900

OpenStudy (anonymous):

Let L and h be lien and her husband respectively. {L=h+1200,L+h=42900} {h = 20850, L = 22050}